Compute C7,5. (Enter an exact number.)
Answer #1

here we have to compute C7,5 means

Combination (7,5)

\\ \binom{7}{5}=\frac{7!}{(7-5)!5!} = \frac{7*6*5!}{2!*5!} \\ \\ =\frac{7*6}{2*1}=21

Hence

C7,5 = 21